40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 13 23:59:46 2010, Zach, Summary, IOO, MC table leveled 
    Reply  Fri May 14 03:30:45 2010, Koji, Summary, IOO, MC table leveled 
       Reply  Fri May 14 10:33:01 2010, Zach, Summary, IOO, MC table leveled 
          Reply  Sat May 15 04:13:33 2010, Koji, Summary, IOO, MC table leveled 
Message ID: 2928     Entry time: Thu May 13 23:59:46 2010     Reply to this: 2929
Author: Zach 
Type: Summary 
Category: IOO 
Subject: MC table leveled 

 After the recent removal of the old IMMT and the relocation of the Faraday isolator, the MC table was tilted a bit (southward and slightly westward---as of when I opened the chamber this afternoon). I re-leveled it by putting an extra two rectangular ballast blocks on the stack that was already hanging off the NNE edge of the table (there are a total of 4 in the stack now). I also screwed down the circular block that Koji and I put between the Faraday and SM1 on Tuesday, and re-mounted the two wire harness towers onto the table.

Needless to say, this threw the MC way out of alignment. I spent the rest of the afternoon reacquiring alignment and getting it to lock robustly. Here is a summary:

  • I adjusted MC3 until I got the 2nd, 3rd+ pass beams to overlap with the input beam between MC1&3, then I adjusted MC2&1 semi-methodically until I got something flashing at the transmitted end. This took some time.
  • I went back into the control room, engaged the loops and acquired lock on the TEM00 mode, whereupon I found that the beam spot was WAY off center on MC2 (due to my meddling with all the mirrors to get resonance flashes). I began using the MC2_spot_up (etc) scripts we wrote the other day to re-center it.
  • After a few iterations, the lock became weak, and eventually gave out. This is because the REFL beam was falling off the RFPD (and being clipped by the iris on the AP table), so I moved the iris and re-centered the beam on the diode.
  • With that, I was able to get the MC2 spot more or less centered, but then I noticed that---though the lock was clearly strong as evidenced both by the REFL power dip and visually via the camera on MC2---it looked like crap on the CCD. It seemed like there was some higher order mode structure sloshing around on top of the 00 spot, which didn't make any sense, until I realized that it was just a diffraction pattern from the TRANS beam getting clipped somewhere on the way out of the vacuum system.
  • I went back to the AP table, where I noticed that the TRANS beam was hitting near the edges of several of the mirrors on the way back to the PSL table, including the first one out of the viewport, so I turned IM4 to center the beam on this mirror, then proceeded to center the beam on each mirror downstream and then onto the CCD.
  • After getting a clear picture of the transmission on the CCD, centering the spot even better on MC2, then fine-tuning MC2&3 to strengthen the lock, I went back to the MC table to check that the transmitted beam was still passing through the center of the Faraday, which, by none other than an act of God, it was.
  • Having done the necessary work in the tank, I ran the A2L_MC2 script to fine-tune the centering of the spot on MC2. It needed a couple steps up and to the side, but after that the actuator gains for pitch and yaw were both balanced again to within ~2%, which is only slightly above the measurement error. We will probably need to adjust this continually, especially during the upgrade, so I didn't bother with getting it better than that.

After that, I shut off the loops, blocked the beam, and put the light doors back on the tanks. Then I went to the parking lot, then I got in my car, etc, etc, etc.

ELOG V3.1.3-